OpenStudy (anonymous):

There are 63 students in a middle school chorus. There are 11 more boys than girls how many boys x and girls y are in the chorus? There are 63 students in a middle school chorus. There are 11 more boys than girls how many boys x and girls y are in the chorus? @Mathematics

OpenStudy (anonymous):

x = y + 11 y + y + 11 = 63 2y + 11 = 63 2y = 63 - 11 = 52 y = 52/2 = 26 x = 37

OpenStudy (anonymous):

M-E-D-A-L

OpenStudy (anonymous):

T-H-A-N-K-S

OpenStudy (anonymous):

can you solve my other one for another medal